Label the points O(0, 0), A(22, 4), and B(1, 1) on the graph
of f (x) 5 x 2. Determine the images of these points on the transformed function. Label the images Or, Ar, and Br.

f(x)=1/2x^2+2
1
Expert's answer
2020-12-06T10:28:07-0500

This question isn't a general chemistry question, it is a precalculus question (linear transformation), but I'll still try what I can do.


Your question isn't clear, but I'll assume you meant

f(x) = 5x + 2

f(x) = 1/2x² + 2

Assume y = f(x)

5x + 2 = 1/2x² + 2

5x = 1/2x²

10x = x²

x = 10

Putting x in 5x + 2 = y

y = 5(10) + 2 = 52


Therefore, point r = (10,52)


With that you should be able to calculate your or, ar or br.